1

Academics

This chapter attempts to provide answers to a variety of questions dealing with

4

Institute academics. For more information, use the cross-reference to MIT Publications in the Publications chapter. The most helpful publications include the Class of 1998 Handbook, the MIT Course Catalogue, the Guide for Undergraduates and Faculty Advisors, and the Graduate School Manual. Entries in this chapter are in alphabetical order.

Advanced Standing

Getting credit for course material you already know is not difficult. For many 12

12

subjects all that is needed is a passing grade on an Advanced Standing exam; for others, additional work may be required. See the instructor and your advisor for details. Incoming students should contact the Admissions Office (3-108, x3-4791). Advanced Standing exams are offered in February, May, September, and December.

22

You must file a petition with the Registrar three weeks before the exam period begins and get the course instructor’s approval. Freshmen are graded pass/no credit; upperclassmen are given letter grades. A freshman failing an Advanced Standing exam will not have a fail recorded on the permanent transcript. Advanced Standing exam grades are never computed into your cumulative average. If you are considering advance-placing a course, don’t register for it—not even as

30

a listener. Registration in a subject automatically makes you ineligible to take the Advanced Standing exam, regardless of whether you remain registered for it or actually attend classes for the course. If you accidentally do this, a petition will be necessary to resolve the problem. Grad students seeking advanced-placement credit generally make informal arrange-

37

ments with the instructor and department rather than take an exam. Most departments allow you to use some subjects taken at other schools to fulfill graduate degree course requirements or the minor requirement.

Advisors

Freshman Advisors can help you with many of your problems during your first year 43

43

at MIT. Get to know your advisor. Associate Advisors are upperclassmen who have volunteered their time to help advise freshmen in conjunction with freshman advisors. If you’re interested in becoming an associate advisor, talk to someone in the UAAO. All freshmen who designate a major at the end of the first year will then be as-

53

signed an advisor in that department. If you know a particular faculty member in your department, you can request him or her as an advisor. Students who do not designate a major before their sophomore year will be assigned a faculty advisor by the Undergraduate Academic Affairs Office (7-104). Your faculty advisor can be very helpful. Remember: She is your link to your

60

department and to the faculty committees (CAP, etc.). If you are not doing well, her intervention can be very helpful. Keep in touch, especially since many advisors are busy with their own work. It is possible to use your advisor as a rubber-stamp signature on various forms, but you’ll have a much better and easier time if you take the initiative and actually discuss your plans and problems with your advisor on a regular basis. This will also qualify her to write your grad-school recommendations.

2 Academics

Changing Advisors. If you’re having trouble with your advisor or have met an-

70

other member of the faculty you would like to have as your advisor, you can change advisors. It is usually not too difficult. Freshmen and undesignated sophomores should talk to the people in the Undergraduate Academic Affairs Office (7-104). They are very helpful. Upperclass students with majors should speak to their department undergraduate headquarters.

Co-op, with industry

Most engineering majors have organized cooperative study plans with industry, some 79

79

of which allow you to complete both your bachelor’s and master’s degree together. Check with your advisor or look in the Course Bulletin for details. The tuition differs from the regular tuition in some departments. Check at department headquarters and the Registrar’s Office for more information.

Committees

Faculty committees make many decisions on academic policy. Committee member-88

88

ship lists are available in the MIT Directory under “Standing Committees of the Faculty,” and can be obtained in 4-237.

1. Committee on Academic Performance (CAP) handles most petitions, in-96

cluding those concerning pass/fail, dropping a course after the drop date, and deadlines for finishing incompletes. They also decide whether to place a student on academic probation. (7-104, x3-4164)

2. Faculty Policy Committee (FPC) maintains a broad overview of the Insti-106

tute’s academic programs, deals with a wide range of policy issues of concern to the Faculty, and coordinates the work of Faculty committees. (4-237, x3-1706)

3. Committee on the Undergraduate Program (CUP) is responsible for en-115

couraging experimental innovation in undergraduate education and formulating proposals for changes in undergraduate educational policy. It exercises oversight responsibility for undergraduate education, including the freshman year and interdepartmental programs, giving special attention to long-term directions. The Committee exercises authority to approve and supervise limited educational experiments.

4. Committee on the Graduate School Policy (CGSP) is concerned with 128

academic performance, degree candidacies, fellowships, and scholarships, etc. It is composed of students and representatives of the graduate committee from each department and program. See your department representative or the Office of the Dean of the Graduate School (3-138, x3-4860) for more information.

5. Committee on Curricula (COC) handles requests for exemptions and sub-138

stitutions in general Institute requirements. It also handles proposals for changing requirements. (E19-341, x3-4781) These committees are always looking for potential active student members. Al-

142

though these committees require a very serious commitment, being on one of them is one of the best ways of leaving your mark on MIT. Student members are nominated by the graduate and undergraduate student governments. For more information call the GSC (x3-2195) or the UA (x3-2696).

Complaints/Suggestions

Your advisor is the first person to turn to. Other natural choices are the department 149

149

head and other professors within the department. The department’s feedback committee, if it exists, can also prove useful. Other resources include the local Course

Co-op, with industry Domestic Year Away 3

society, the Executive Officer, or even the secretaries at Headquarters. (These last offer information rather than brute-force pull. However, that information may prove invaluable and is unavailable elsewhere.) The UAAO (7-104, x3-6771) has no formal departmental connections but can give good advice.

Cross-Registration

Sometimes a course that you want to take is not offered at MIT. But not to worry— 163

163

cross-registration programs with Harvard and Wellesley open the door to many fields not available at the Institute. Harvard. Cross-registration at Harvard is now open to all MIT students. Fresh-

169

men, however, are restricted to foreign language courses not offered at MIT (such as Arabic or Korean). Generally, you cannot take an MIT-equivalent subject at Harvard unless you have a serious schedule conflict between the MIT version of the course and another MIT course. See Bette Davis (14N-408, x3-4443) for Harvard catalog listings and other information. You can also check with the Harvard University Information Center at Holyoke Center in Harvard Square. Harvard’s deadline for registration falls shortly after MIT’s registration day, and the

179

calendar is different from MIT’s; first term finals are in January, and second term finals are in late May, often after MIT’s finals. Be sure you will be around during their finals week. Wellesley. In the Wellesley Exchange office (7-103, x3-1668) you will find infor-

185

mation sheets on cross-registration, bus schedules, Wellesley catalogs, course evaluations by MIT students, and more. Information about the exchange is also posted on the exchange bulletin board near the Admissions Office. Wellesley classes begin about a week before MIT classes do, so it is important to

192

begin attending Wellesley classes no later than the first week of MIT classes. You must cross-register at Wellesley during the first week of classes. MIT needs you to list the Wellesley subject on your original registration form or fill out an add card. A free bus service is run between MIT and Wellesley Monday through Friday after-

198

noon. The cross registration bus schedule is based on a combination of MIT and Wellesley class schedules. Be sure to allow enough time for transportation between classes. Note that some Wellesley classes are taught at MIT. Policies concerning use of Wellesley subjects to meet general Institute and depart-

204

mental requirements are described on the information sheets. Information about course prerequisites, the meaning of “limited enrollment,” or anything else can be gotten from the Wellesley instructor. For more information see Wellesley in Colleges Around Boston.

210

Degree requirements

See Graduation and Degrees.

213

213

Domestic Year Away

You can spend a year studying at another college in the United States provided 216

216

that certain conditions are met. The school you want to attend must accept you and be of “established merit,” and your department must certify that the school has unique resources unavailable at MIT, Wellesley, or Harvard. Your program of study there should involve a workload comparable to MIT’s, and its objectives must be consistent with your overall MIT degree program. You won’t have to reapply to MIT to get back in.

Financial aid is available for those who would normally be eligible. The Student

229

Financial Aid Office will determine your need at the other school and what your need would have been at MIT and will give you the smaller award. You can get more information from the Foreign Study Office within the Career Services office (12-185, x3-4733).

End-of-Term Regulations

In a course with a final exam, no written exams may be given during the six days 235

235

before the reading period before final exams. Also, no assignments or term papers may be due. In courses without finals, either one in-class exam or one assignment may be due during the six days precding the reading period. No term papers or assignments are allowed to have due dates after the last scheduled class meeting of any class. It is not uncommon for professors, especially those new to MIT, to be unaware of

246

this policy or to accidentally violate it. (It also happens, but is rare, that a professor will intentionally violate the policy.) If you think there is a problem, first talk to the professor (possibly with other members of the class). If you are not satisfied with the result, call the UAAO (x3-4164) to complain to the chairman of the faculty.

Finals

Final exam schedules come out in midterm, and are available at the UAAO and the 254

254

Information Center (7-121). If you have conflicting finals, follow the instructions on the Conflict Form on the back of the schedule. (See the section on End-of-Term Regulations to see what is permissible during the last week of a class with a final.) Call the UAAO (7-104, x3-6771) if you forgot (or never found out) where your final

264

is supposed to be. The Schedules Office (E19-339, x3-4788) has the most up-to-date information, and they handle conflicts. If you will be more than 45 minutes late, you must get permission from the SAS (Student Assistance Services) office (5-106, x3-4861) to enter the exam room. If you cannot take a final at the scheduled time, all is not lost. Undergrads should

271

go to the SAS Office (beforehand if possible) and explain the circumstances which prevent you from taking the exam (a personal visit is strongly preferred to a call). If your reason is accepted, and the instructor has issued an “O”(absent from final or last two weeks of class up to that time), the Dean’s Office will issue a grade of OX. It is your responsibility to make up that final. If you had been doing failing work throughout much of the term, the instructor still has the right to give you an F instead of an O. If you are sick, be sure to go to the Medical Department before the exam. If your

281

reason is not accepted, the O remains as such on your record—it is equivalent to an F and is averaged into your GPA. A postponed final is normally taken early in the next semester. This means that if

285

you missed a second-term final, you must retain all of your knowledge of the course through the summer and part of the fall term. So it might be better to take the final on time if you are not too sick. Be realistic about your capabilities though, because if you do badly, it is nearly impossible to have the grade changed. Graduate students who need to reschedule a final should talk with their instructor. 292 Also, you can check with the Graduate Office (3-138, x3-4860). If you have a tendency to go to pieces or otherwise fail to perform maximally on

296

End-of-Term Regulations Freshman Troubles 5

finals, talk with your instructor beforehand; afterwards pleading for an incomplete may be your best option.

Foreign Study

If you are interested in studying abroad you should start to plan as soon as possible. 300

300

Junior Year Abroad is generally the easiest way to go. Student status and dorm priority are generally unaffected. Financial aid is available as for Domestic Year Away. Contact the Career Services Office (12-185, x3-4733) for more information. The opportunities for foreign graduate study are greater. The Office of Career

312

Services and Preprofessional Advising (12-170) has some reference materials. The Graduate School Office (3-138, x3-4860) also has information on DAAD, Churchill, Fulbright-Hays, Marshall, and other foreign scholarships.

Freshman Troubles

Not everyone’s freshman year is a non-stop joyride. You may encounter one or more 318

318

of the following difficulties: Advisor problems. If you don’t like your advisor, you can ask to change advisors

321

at the Undergraduate Academic Affairs Office (7-104, x3-6771). Tell the UAAO people what you want in an advisor. Don’t hesitate to request a change if you feel one is necessary. You can change advisors more than once as well. (See Advisors above.) Evaluation forms. If you haven’t gotten them in the mail, you can get them

328

from the UAAO (7-104). If you don’t get the completed forms back, bother your instructor. These forms may feel silly, but they help give you a good gauge of how you’re doing. Failure (not passing). Don’t panic. Normally you will merely have to repeat

334

the course if you want to get credit for it. Pleading sometimes helps. Possibly your evaluation forms did not give you a clear idea of where you stood. Talk to your professor. Remember that freshman failing grades are not recorded on your external transcript (but they will influence the CAP at the end of the term if you fail enough classes to be put on academic probation). Seminar problems. If you don’t like—or think you won’t do well in—your sem-

342

inar, you can drop it. However, since most seminars are oversubscribed, give your seminar a chance. You probably won’t be able to switch to another. However much these seminars have been advertised as “fun” courses, they are sometimes treated as an indicator of your creativity and research potential. Beware that a few “freshman” seminars have, on occasion, assumed background beyond that of the average freshman. (One year, a number-theory seminar assumed substantial familiarity with group theory.) If that’s the case, talk to your seminar leader. Probably everyone else is floudering, too. You can obtain a listing of the seminars for the next term at the Undergraduate Acacdemic Support Office in 7-104 (x3-6771). Course Problems. If you really don’t like your professor, go to the undergraduate

357

department office or to the departmental headquarters and explain why you would rather be in another section. Don’t use the negative approach; instead, explain why another section would be better for you. If there is only one section, you can drop the course (and wait for next term) or grin and bear it. If you have any problems at all, whether they concern problem set difficulties or

366

room temperature, talk to the instructor. She wants feedback (in most cases) and will react favorably to the simple fact that you have gotten off your tail and said something—even if it is critical. Just do things tactfully, and you will be surprised at how well received you’ll be. If that doesn’t work, try your advisor or other faculty members in the appropriate

373

department. Go straight to the department head if necessary. Try the deans, especially the UAAO (7-104, 3-6771), and then the counseling staff (5-106, x3-4861); they are all experienced in dealing with classroom complaints. Don’t worry that your comments might be too trivial to waste someone’s time with.

379

If something bothers you, it probably bothers other people, too. If you’re really unsure, sound out another student, a dean, or a random faculty member and see what happens. Normally, all you have to do is ask. If your problems involve the course material, you have several paths available to

385

you. Tutorials (if they exist) provide personalized help—it’s best to attend them throughout the term and avoid the crowded cram sessions on the day before the exam. Your TA might also be able to arrange a help session. Faculty and students in your living group may have taken the course in question; they might have old quizzes and homework as well as an understanding of the material. Undergraduate offices also often have facilities to help out. The Office of Minority Education (7-143, x3-5010) offers tutoring for all students. Schedule Conflicts. If you have two or more conflicting subjects scheduled,

396

check with the Schedules Office (E19-338, x3-4788). If you can’t get into a different section, you’ll have to go to half of the classes in each subject. Dropping one of the courses and waiting until the next term may be the best solution. Discuss the problem with each one of the instuctors. The schedule given on your schedule card is taken from the master list given in

407

your registration booklet. During the first week of classes, many courses will not pay attention to the section assignments and you can simply show up at the most convenient section. However, this is not true for all courses; many have special procedures to change sections. If you’re not sure, it’s generally best to ask the instructor of the section you want to attend what you should do to remain in the section. After the first week, you may need departmental approval to change your section.

Grade Challenging

Problem sets, quizzes, and exams are often marked by TA’s or graders. If you have 417

417

any questions about the accuracy of a mark, ask the grader or TA; if you cannot resolve the problem, talk with the instructor. Normally he will be quite willing to look at your side of it. If you feel a course grade is unfair, first talk to the instructor, and then, if necessary,

424

to the department head. The CAP is willing to act as an intermediary in case of disagreements but cannot override a department’s decision. The instructor must fill out a Special Report Sheet and send it to the Registrar before the grade change can officially take place. All such changes, including making up I’s and OX’s, must be done before you get your degree.

Graduate Schools

Reasonably thorough collections of graduate school catalogs can be found in the 432

432

Humanities Library (in 14S) and the Career Planning and Placement Office (12-170, x3-4733). The latter also has independent references on school reputations,

Grade Challenging Graduate cross-registration 7

financial aid, and degrees awarded. In addition, the office has some applications for the GMAT, LSAT, MCAT and GRE. With regard to the GRE, the Graduate School office (3-138, x3-4860) has application forms and information. Reference copies of test bulletins are kept in the Career Office (12-170). Register to take the GRE’s and other such exams spring term your Junior year. Frank Perkins, Dean of the Graduate School, can be helpful to talk to if you want

447

to come to MIT for graduate study. When applying to grad schools, it’s usually best to specify the doctoral degree as the “degree to be studied for.” This and other rules of thumb may be gleaned from your advisor and other faculty if you ask for help. Getting A Recommendation. Recommendations may be at least as important

453

as your average. You should have a good working relationship with at least one faculty member and ask for recommendations early. Make sure that you ask people for their approval before using them as a reference. It helps to know that you’ll get a good recommendation. (Unlike many people in the “real world,” MIT people tend to be brutally honest in their recommendations.) Admission to MIT Graduate School. For some departments (particularly

463

Chemistry and Biology) it will be harder for MIT undergrads to get into MIT than for grads from other institutions of comparable quality. The reason is that departments want to avoid “inbreeding.” Consult the department, the Admissions Office, and the Graduate School Manual for more information and help. It is possible to be admitted as an interdisciplinary graduate student. You still have

473

to find a department that will admit you, but some flexibility is allowed. It will take a lot of talking to get you what you want and at least one professional ally. Don’t delay. Prelaw, Premed, and Educational Counseling. The Committee on Prepro-

478

fessional Advising and Education serves the special needs of students interested in entering the medical, legal, and teaching professions (and related areas in public administration). The Committee itself consists of three advisory councils, one in each field. If you are wondering what it feels like to be a professional in one of these areas, the members of the advisory councils will be glad to tell you. The Office of Career Services and Preprofessional Advising (12-170, x3-4737) pub-

492

lishes special handbooks for the fields mentioned above.It also provides graduate school catalogs, applications for entrance exams (MCAT, LSAT, etc.), and academic counseling and sponsors seminars given by professionals and representatives of graduate schools. Students interested in teaching should contact the Wellesley Education Department. MIT students can gain teaching certification through Wellesley. See the Exchange Office for details (7-103, x3-1668).

Graduate cross-registration

MIT has cross registration programs with a number of schools. Most of these pro-505

505

grams are very limited (e.g., at Brandeis only course XI grad students may cross register), but don’t let these restrictions stop you. Build a good case for yourself, convince your advisor of its merit, and you at least have a chance. With perserverence and careful planning, things can go the way you want them to. More information is available in the Graduate Student Manual and from your registration officer, or from the appropriate coordinator:

School . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . MIT Coordinator Harvard . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . Registrar’s Office, E19-339 Wellesley . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . Wellesley Exchange Office, 7-108, x3-1668 Woods Hole Oceanographic Biology; Ocean Eng.; Earth,

Institution . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . Atmospheric and Planetary Sciences Brandeis (Social Welfare) . . . . . . . . . . . . . . . . . . . Prof. John Howard, x3-7333, x3-4408 Tufts (Dental School) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . Departmant of Applied Biology Boston University (African Studies) . . . . . . . . . . . . . . Economics and Political Science

Graduation and Degrees

You must submit an application to be a degree candidate in order to graduate. The 534

534

application is included in your registration material. MIT does not automatically grant your degree as soon as all your requirements have been met. It is possible to graduate with one deficiency if it occurred in a departmental subject for which you were registered in the final term and you made an honest effort to complete the course. Even then, graduation is not automatic. Your department will have to petition the Committee on Academic Performance (CAP) and make a well-informed plea. Double-majors may not graduate with a deficiency. Before you receive your diploma, all library books must be returned, all keys to

546

rooms and labs must be returned to the Office of the Superintendent of Physical Plant or to the appropriate professor, and all financial matters must be cleared. The Student Accounts office will notify you of the deadline dates for all the formalities. Prior to graduation, any student employed by the Institute must file a Termination

552

Clearance form with the Payroll Office, with signatures certifying that she has returned all keys, books, supplies, etc. Some labs have their own special termination procedures for teaching assistants. Two Bachelors Degrees: You must fulfill requirements in two departments while

562

accumulating an additional 90 units (about one year’s worth). You may petition the Committee on Curricula (through the Registrar’s Office) for a double major only after completing two semesters on grades. They say that you must petition at least two terms before graduating; but they have been known to relax this requirement if you have been making a demonstrable effort towards the second degree. The petition must list the two degrees and the expected graduation date. Consult your advisors in both departments to find out about the deadline for application and to obtain approval of your entire program. Make sure you contact both departments about your joint thesis if you plan to

574

present one. Bachelor’s and Master’s Degrees: Some departments will award an SB and an

577

SM simultaneously. You must be accepted for graduate study at least one term before the degree is awarded; that requires a petition. Check with your advisor for details. If you want a Bachelor’s degree in one department and a Master’s in another, you are on your own. Don’t worry, though—it has been done. Two Master’s Degrees: You do not necessarily have to take additional credit

585

units for two Master’s degrees. However, there are requirements about splitting up the units between the two departments. You can do a joint thesis for two Master’s degrees; a petition must be filed at least two terms before graduation. Degrees, Interdepartmental. Many departments have concentrations which

592

allow much work in another area (II-A, VI-1 Bio option, etc.). Check with your advisor.

Graduation and Degrees Incompletes 9

Degrees, Unspecified. Several courses (I, II, III, IV, VII, IX, X, XII, XIII, XV)

596

sponsor very flexible programs that can be adjusted to suit your personal preferences up to a certain limit. Since all departments have the option of offering unspecified degrees, your advisor or department head can help you set up a program. But remember that you still must fulfill the general Institute requirements, even if they don’t seem to fit into your personalized educational program. How to Get Around Requirements. If you want to be exempt from making

606

substitutions for Institute requirements, you have to petition the COC. Getting out of a departmental requirement will require working things out with your advisor and the departmental head. Talk to the people involved before writing your petition. Your advisor and the deans can be of help. Try to find out about policies and precedents, and build up a believable case, which may take a major creative effort in borderline cases. Don’t be afraid to try; the results can be worth it. Petition forms are available

619

from your advisor, the Registrar’s Office, the UAAO, and the Information Office. Your petition will be more effective if you justify yourself in two sentences on the petition itself. Items which are petitioned for most frequently are substitutions for humanities and lab requirements. Often other subjects may be substituted for certain requirements in the department. Make sure you have the approval of the instructor in the subject you want to substitute. If your petition is refused, you can appeal to the same committee and they may change their minds if you are persuasive enough.

IAP

IAP is a unique MIT experience. It is the time in between fall and spring terms— 630

630

about 4 weeks in January. During this time there are many activites, seminars, and yes, even a few accelerated courses. It is a time to explore some of the things which have always interested you but that you haven’t had time for. If you have some skill or knowledge that you would like to share with others, you can teach your own course or sponsor your own activity. Past topics taught by professors and students alike include wine tasting, machining, algebraic topology, sightseeing, and a real-time roleplaying game. There is no real limit to what you can do. Information about organization, registration, and funding (it exists!) is available in

644

October from the stacks of forms in Lobby 7 or the IAP office (7-103, x3-1668). If you have any questions, ask Alice LaPierre, who is very friendly and helpful. The IAP Preview, published in September, contains major activities planned by academic departments, but a complete list of the 600-plus IAP activities is not available until early December in the IAP Guide. Credit for IAP. You are generally limited to twelve units of credit for IAP. Some

652

12- and 6-unit classes which meet departmental requirements for some majors are available during IAP. Taking over twelve units over IAP requires a petition to the COC or CGSP. The freshman credit limit does not apply to IAP credit.

Incompletes

Incompletes must be completed by the end of the fifth week of the succeeding term, 659

659

unless the instructor explicitly grants you an extension (which at most extends to the last day of classes of that term). Any further extension requires the approval of the CAP, which is given only in special circumstances (for instance, if the lab equipment you need is not available during the appropriate term). If you withdraw from MIT before making up the incomplete, talk to the CAP about when to make it up; if the incomplete requires scarce MIT equipment, remember that you are not officially allowed to use such equipment when not registered. The CAP encourages readmitted students to make up incompletes. Incompletes not completed remain Incompletes; they are not computed into your

672

average. However, they are often treated by graduate schools as “F”s. Departments and the CAP will consider the number of incompletes in your record if your academic performance is being reviewed. It is a good idea to ask your professor ahead of time for an Incomplete (but don’t

678

ask if you haven’t already done two thirds of the course work), explain the reasons for it, and tell him when you plan to finish the subject. (If he is not going to give you an Incomplete, better to know before the term ends.) Many professors will go out of their way to be accommodating, even to the point of calling you up to see if you are still working on the Incomplete; others will screw you to the wall.

Inventing Your Own Course

New courses have been invented in the past. A strong commitment and faculty 687

687

backing are essential. The Educational Studies Program (ESP, W20-467, x3-4882) provides an organization for teaching Saturday courses to high school students. Starting an IAP course is much easier and can be done through a department or independently. Contact the IAP office (7-103, x3-1668).

Light Load

Regular students planning to register for a light load (fewer than 33 units) must get 696

696

approval from their advisors and a counseling dean (5-106). If the forms are signed by add date, the CAP will take this into account at its end-of-term meeting and will not take action because of a low registration. There is also the possibility of saving money by paying tuition on a per-unit basis while still being considered a regular student. Taking more than two terms of light load is discouraged and requires a CAP petition.

Listener Status

There are two ways to sit in on a subject: the first is getting permission to sit in on 706

706

a class and learn informally; the second is registering officially to listen to a class. If you do register to listen, you will not be allowed to advance-place the course later on. For the summer term, or whenever you pay tuition on a per-unit basis, the listening rates are the same as for regular students’ tuition. It is possible to change status to a regular student during the term (up to the add date). See your advisor for the necessary paperwork.

Pass/Fail

Pass/Fail grading is generally available only to freshmen, to students in seminars 716

716

or UROP programs, and to juniors and seniors exercising their pass-fail option (see below). An instructor may petition the COC to have his/her entire subject made pass/fail but cannot allow individual students to take the course pass/fail unless one of the above options applies to them. Freshmen will receive only pass/D/F grading on their grade reports, but their hidden

724

grades from the second term will be sent to their sophomore advisors on a separate sheet. (Be aware that the Math, Physics, and Chemistry departments keep records of how first-term freshmen do on quizzes, problem sets, etc.) Your external transcript will not show “no-record” grades, or D or F.

Inventing Your Own Course Registration and Course Selection 11

The junior-senior option allows you to designate two subjects in your last four terms

734

as pass/fail. The subjects cannot be Institute or departmental requirements, but you may take both in the same term. You can change a class to or from pass/fail until add date. Failing grades under junior/senior pass/fail are not recorded.

Problems

If something is wrong and you can’t find adequate help from your advisor, the 740

740

UAAO, the deans in 5-106 (x3-4891), or anywhere else, go to the psychiatrists (E23-376, x3-2916). These people specialize in helping when you’re in trouble.

Registration and Course Selection

The first time you register you will be given registration material by your advisor. 746

746

After that, you pick it up in Lobby 10 on certain pre-announced dates. Spring term registration forms are given out in early December, fall term forms in early May. If you miss the scheduled dates, you can get the material at the Registrar’s Office (E19-335). You must return the forms on time, even if the information is incomplete; otherwise,

756

you’ll be stuck with a $20.00 fine. You can always make changes on Registration Day. (If you really have no idea what to sign up for, just register for anything and change it later.) The Registrar will send you notification of your registration status periodically

762

throughout the term (more-or-less after Reg Day and before and after Add and Drop Dates). Check these carefully and be sure to have them corrected immediately if necessary. Be sure to keep a copy of everything you can whenever you deal with these people, especially end-of-term grade reports. The Registrar’s computer has been known to “forget” you registered for a 24-unit lab, etc. Don’t count on the experience of others to prepare you for what may come—this office is also very innovative. To add or drop a subject before the applicable deadline, submit a correction card

772

(“add/drop card”) to the Registrar’s Office (E19-335) or the Registrar’s Annex Office (7-108). Cards are available at the Undergraduate Academic Affairs Office (7-104), the Information Office (7-121), and some departmental offices; deans and advisors can sometimes supply them. The card needs your advisor’s signature. If you are adding a course after the first week, you also need the instructor’s signature. If your advisor is not available, a signature from your department office (or the UAAO for freshmen and undesignated sophomores) may be obtained. A course dropped within the first five weeks of the term will be entirely erased from your records; a drop after five weeks will appear on the internal transcript. Keep in mind that some faculty members consider a drop after the fifth week to be

788

automatic proof that you were failing the course at the time. After Drop Date (eleven weeks from the start of the term), you must petition the

792

Committee on Academic Performance if you want to drop a subject. The CAP has outlined adequate reasons for dropping a subject late as unforeseen circumstances beyond the student’s control. The fact that you are failing, by itself, is not enough to get you off the hook. As a whole, the committee is very strict about drop date, but exceptions have been made in the past; check with the staff working with the CAP (7-104, x3-4164) to see what your chances are. If you have an exam on drop date, don’t wait until the next week to see your score

801

and then petition to drop the course. If you have an exam the day before drop date, you can often obtain your scores early by calling the professor or a TA for the course. If you’re not sure whether to drop or not, visit your professor or recitation instructor during office hours or make an appointment with them. They will tell you whether you should stick with it, get an Incomplete, or drop now while you still can.

Special Projects

If you have an idea you want to work on or something you want to study, it is 810

810

possible to get Institute backing. You may even be able to get credit or satisfy Institute requirements while doing your own thing. Petitions for fulfillment of Institute requirements go to the COC. First, dig up an idea. (Example: writing a guidebook for the MIT community;

817

that’s how this book got started.) Several established ways of doing so are:

1. Check with the UROP office (20B-141, x3-7306) or check their bulletin board in 821

the main corridor for current research offerings. Although most projects take place on campus, it is also possible to have an off-campus project at hospitals, corporations, non-profit organizations, or government agencies. See the UROP booklet for more ideas and for the rules and regulations regarding project work. Also, see the UROP section of this chapter.

2. Ask at department headquarters/graduate offices for references on current re-829

search.

3. Find a professor you want to work with and ask her for project ideas. Most 832

faculty members have lots of ideas which they themselves don’t have time to work on.

4. Drop in at a lab you’re interested in and ask someone about what’s going on. In 836

most cases, you’ll get an enormously thorough description; people are very willing to talk about their work. Second, find a way to do your own thing. A few suggestions are:

840

1. Register for a “projects” or “special problems” subject in your (or any other) 842

department. The best types are those with credit “to be arranged.” You must find a faculty sponsor. If the project works out well, you can do a careful write-up and petition to make it your thesis.

2. Try to make your project a part-time job with a professor or try to make it 848

a summer job, possibly by getting hold of grant money. Consult your advisor or department head for suggestions on how to go about it.

3. Check with the interdepartmental labs. A fairly thorough listing is in the Catalog. 853

Whatever your idea, if you’re really interested in pursuing it you should be able to

856

find help and sponsorship somewhere in the Institute. If at first you don’t succeed, keep trying; there are plenty of places to go for advice. The person who turned you down may change his mind and offer his services the next month.

Special Students

Special students, by definition, are not considered to be working toward a degree. 862

862

Special students pay tuition based upon the number of units they register for; they must be readmitted each term. They usually are not eligible for financial aid, campus housing, or cross-registration. Subjects taken by a special student can be used toward a degree if the student is subsequently admitted as a regular student. Contact

Special Projects Transferring Credit 13

the counseling deans (5-106, x3-4891) if you’re thinking of changing your status or the Admissions Office if you’ve never been a regular student.

Summer Session

The summer session is open to all students. The summer catalogue is issued in 873

873

March and is available at the Information Office (7-121) and at the UAAO (7-104). Tuition is paid on a per-unit basis although minimum and maximum rates do exist. Regular students should check with the Registrar’s Office (E19-335); Special Students should contact the Admissions Office (3-108). Non-MIT students cannot register for MIT summer session without being admitted

881

to the Institute under the same admission standards as those for the regular school year. The number of courses offered in the summer is limited, but it usually includes some

885

large freshman and sophomore courses such as 8.02 and 18.03. Summer is a good time to do research. It can be done for credit or pay. If you do

890

it for credit, however, you will have to pay for it, depending on how many units you register for. The research must be started at least halfway through the spring term, especially if you want to be paid for it. If you want to do research for pay, funding either from your department or UROP may be difficult to obtain. Check with the UROP Office (20B-141, x3-7306) for more information and help, and check with your department about ongoing projects which you might find attractive.

Theses

Copies of Specifications for Thesis Preparation, published by the Libraries, are avail-900

900

able from your department, the UAAO (7-104), and the Institute Archives (14N-118). Your advisor or department should be able to answer most questions.

Transcripts

The Registrar’s Office will make copies of your transcript for $2.00 each. Keep 906

906

in mind that it will take a while to get them prepared. Payment is made at the Cashier’s Office (10-180). Requests for transcripts must be made by the student in writing; no request from a graduate school or anyone else is honored without the student’s approval.

Transferring Credit

It is possible to receive MIT credit for work done at other institutions. If an exact 914

914

MIT equivalent exists, fill out a Request for Additional Credit form with a recorded grade of “S.” If you want a real letter grade, consider taking an advance-standing test rather than transferring the credit. (Make sure you look at the MIT problem sets and exams first!) There are transfer credit examiners in each department and Humanities sections.

923

The UAAO has a list of them. You should check with them before taking a course to make sure you will get the credit you want. If no equivalent exists here, see the most closely related MIT department. Any questions about the granting of credit for work done before you entered MIT as a freshman should be directed to the Director of Advanced Placement in the Admissions Office. Graduate level subjects completed satisfactorily at other universities may be ac-

931

cepted toward requirements for an advanced degree but do not contribute to the residency requirement. If the subject has an MIT equivalent, submit an Additional Credit sheet; if not, you must petition the COC. Courses previously taken by a Special Student may be used later in partial fulfillment of requirements for a graduate degree. In either case, the Registration Officer should indicate the amount of (A) or other credit accepted.

UROP

The Undergraduate Research Opportunities Program (UROP) is rooted in ideas 940

940

presented in a lecture by Edwin Land in 1959, when he argued that every student should have a faculty guide to help the student explore his/her own connections to learning and inventing. Land, the founder of Polaroid Corporation, used to sneak into New York University labs at night to try out his ideas about the polarization of light, because he didn’t have anything like UROP to help him. These days a lot of students think about UROP simply as a way to get an interesting

949

research job, and the program works quite well for this. Many students find that UROP leads to thesis work, to graduate school, and to careers. In fact, the people at UROP say they will let you do “anything you want, as long as you can get a faculty member to supervise you (and the department to approve).” Thus, while there are hundreds of undergraduates building their research skills while

956

working on projects for high-powered MIT faculty, there are also a goodly number developing their own projects in science, technology, and the humanities. You can start a UROP any time during your four years at MIT. If you are a first-

962

year student, UROP has a program called Student Research Partners, which will match you up with an experienced upper-classman on a UROP project during the Independent Activities Period in January. UROP makes it possible for you to get either academic credit, or pay, for projects

967

done through the program. The bright yellow UROP directory is a useful tool, and so are the bulletin boards in the Infinite Corridor and the Undergraduate Education Office over in building 20. See the staff in 20B-141, x3-7909. They are very friendly and very helpful. Credit for UROP. First, you need to arrange with a faculty supervisor such

974

details as the nature of the project and the number of units to be awarded. Next, you need to compose a Letter of Intent and have it signed by your supervisor and submitted to the UROP coordinator in your faculty supervisor’s department. Register for UROP credit as you would for any class. If you want to receive pass/fail

981

credit, register for the UR subject number (e.g., 1 U.R., 2 U.R.). If you wish to receive a grade, register under the suitable subject number for undergraduate research (e.g., 7.90, 17.901, etc.). A UROP may be added after Add Date if it really starts after Add Date. And don’t forget to remind your faculty supervisor to submit a grade at the end of the term—if your research isn’t finished, a grade of “J” may be awarded to indicate satisfactory research activity continuing beyond one semester.

Undesignated Sophomore Status

The Undergraduate Academic Affairs Office (7-104, x3-6771) works with students 991

991

who have not designated a major. The people there will talk to you about selecting a major and choosing or changing advisors. They can often help with problems of course and career selection. If you choose a major and then change your mind, don’t worry. It is possible for

998

sophomores to un-designate.

Warning

UROP Withdrawal 15

The description of Warning below is quoted from a publication of the Committee 1001

1001

on Academic Performance.

“Warning”: This action would be appropriate in the case of a student whose per-

1006

formance is considered below standard, and whose status at the Institute may be jeopardized if his/her performance does not improve in the coming term. We feel that this action ought not to be viewed as punitive, but rather as a helping mechanism that will encourage such students to plan constructive action to improve their performance.

Undergraduates. If you have a problem or are facing a warning, talk to some of

1014

the Committee members or to the staff of the CAP (7-104, x3-4164). Most of these people are quite reasonable individuals. There is no set cut-off point for getting a warning or escaping it. CAP always will

1020

look at you if your term GPA is 3.0 or less, if you were registered for less than 36 units at the end of that term, or if you have received more than 15 units of Incomplete. The normal procedure for each department is to have a grades meeting at the end of each term, at which they will consider each student and make recommendations to the CAP. Therefore, if you think there is something the faculty of your department should know, tell them via your advisor or favorite professor before the end of the term. Your advisor is your representative and advocate at these meetings and before the

1030

CAP. If you are having problems, make sure that she knows the full situation before you get into trouble. If it’s too late for that, call her as soon as possible to provide your side of the story. If you have been ill, be sure you have thorough medical records to show the CAP,

1036

as you will not be believed otherwise. If you are claiming you have had “emotional problems,” you may have to get verification from a psychiatrist. If you are on warning, your registration is limited to a maximum of 48 units (51

1041

units if you’re taking a 15-unit course). In some instances, the CAP may vote to restrict a person’s unit load to 39 or 36 units. If you leave the Institute for academic reasons, you should consider yourself on

1046

academic warning the first term after you return—the CAP does. If you have any questions, go to the UAAO (7-104, x3-4164). They can also refer

1050

you to a member of the committee. In general, the committee is quite reasonable and tries to help students. Graduate Students. Graduate students may receive a warning from the Com-

1054

mittee on Graduate Student Policy if their cumulative GPA is somewhere below the 3.5-to-4.0 range. However, the Committee is not required to warn you. If your poor performance continues, your department will probably recommend kicking you out—it is vital that both the CGSP and your advisor clearly understand the circumstances which are causing your problems.

Withdrawal

If you have been on warning at some point and are now doing poorly, the Com-1064

1064

mittee on Academic Performance may, after consulting your advisor, require you to withdraw. Virtually no one flunks out due to lack of intelligence; if you’re having trouble, there’s probably some other factor at work.

If you’re headed for a withdrawal, don’t panic. Visit the counseling deans (5-106) or

1071

the psychiatrists (E23-376). Feel free to ask any questions in the CAP office (7-104). The Placement Office (12-170, x3-4733) can help you find a permanent job (or a temporary one if you intend to reapply for admission). If you plan to transfer to another school, remember that most universities require a

1078

minimum 4.0 GPA (on the MIT scale) for transfer students. If yours is substantially below that, don’t be surprised if you are told to go to junior college or night school. Some schools will allow you to “prove yourself” as a special student and then transfer in, but many will only care about your MIT GPA and won’t let you in even if you have five terms of straight A’s as a special student. If you really think you can’t do MIT work, transfer out before the end of your sophomore year; in that case your high school grades will be the major basis for the admissions decision, and your MIT GPA won’t hurt you that much. Northeastern University in Boston has accepted quite a number of people who have flunked out of MIT, but this school does not provide a particularly good undergraduate education in technical areas. Readmission after Required Withdrawal. Normally students seeking read-

1093

mission after one required withdrawal apply through the Counseling Deans (5-106). Students with more than one required withdrawal must apply for readmission through the CAP. The character of the CAP can change quite a bit. Note that the CAP’s membership

1098

changes partially every year and the committee that can readmit you may have very different attitudes than the one that kicked you out. When applying for readmission through the CAP, you should submit every existing

1103

piece of evidence showing you are no longer a “person with problems” and will be a productive, disciplined, and enthusiastic student. This includes job recommendations, transcripts, and a medical report if you think your past problems were medical in origin. See the UAAO (7-104, x3-4164) for help in writing an effective application. If applying for readmission through the Counseling Deans, talk to other students

1111

who have gone through Dean’s Office readmission before selecting any particular dean. (In the past, some deans have tried to restrict a student’s choice of major without much investigation into the student’s strengths and weaknesses. Be aware that many people regard pure mathematics and physics as unsuitable for someone who has had academic problems.) Voluntary Withdrawal and Leave of Absence. If you want to leave school,

1119

even if only temporarily, you should contact the counseling deans (5-106) for advice and help with all the procedural hassles. If you’re not sure whether to take a term or so off, try talking to one of the deans, your advisor, friends, your personal physician, or members of the Psychiatric Service. Also, the Placement Office can help you in finding a temporary or permanent job. When you request a leave of absence, you will be asked to indicate when you plan to return, but you can change your return date later. Readmission of undergraduates who have withdrawn in good standing is not hard

1130

and is done through the DSA. You won’t lose financial aid, and housing spaces are often available. Readmission of graduate students is done through the department. If you’re gone

1134

more than five years you’ll also need the approval of the CGSP.